# MagicColor: Multi-Instance Sketch Colorization

### v1

@@ 5.1 Implementation Detail
Dataset. The dataset employed in this study is comprised of two key components: the anime video dataset and the image dataset. The data were preprocessed from Sakuga [57], ATD-12K [67] and manually collected from the internet. A total of 1,000 image pairs were randomly selected to form the test set and the remaining data constituted the training set. We then extracted their instances as references.
@@
Hyper-parameters. During training, we resized the height and weight of the input image to 512 and kept the original aspect ratio. Our model undergoes 100,000 optimization iterations with a batch size of 1. The learning rate is set to $1\times 10^{-5}$. The training phase takes around 7 days using 2 NVIDIA A800 80G GPUs. We perform a random horizon flip and random brightness adjustment for each reference image as data augmentation to simulate multi-view conditions.
@@ 5.2.1 Qualitative Results
We compare our approach with previous sketch colorization methods. For GAN-based approaches, we consider RSIC [30] and SGA [36]. RSIC and SGA utilize GANs for colorization, each with a distinct architecture and training strategy. We observe that GAN-based methods tend to produce color incoordination. Moreover, when there are numerous color block regions in the image, color bleeding issues are likely to occur, where the color of one color block contaminates the adjacent color blocks. Regarding diffusion-based models, AnimeDiffusion [6], colorizeDiffusion [90], and MangaNinja[45] (without point guided) perform better in terms of color control. However, as shown in Figure 8, they still struggle to effectively transfer the colors from the reference instances to the line drawing sketch, especially for the colors of smaller details in the images. The root cause is that the models fail to fully learn the color correspondences between the original image and the reference instances.
Table 1: Quantitative comparisons with SOTA baselines.
FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
RSIC[36] 78.954 20.527 0.719 0.395
SGA[40] 101.384 18.771 0.679 0.436
AnimeDiffusion[6] 118.85 11.037 0.482 0.549
ColorizeDiffusion[90] 67.161 10.680 0.435 0.515
MangaNinja[45] 84.239 11.608 0.434 0.577
Ours 28.443 23.493 0.805 0.277
Figure 7: User study results. Quality: Visual fidelity of generated outputs, Similarity: Alignment with reference instances, Usability: Operational simplicity, Controllability: Consistency between references and result outputs.
@@ 5.2.2 Quantitative Results
To comprehensively evaluate the colorization ability of our model, we quantitatively compared our method with the state-of-the-art colorization method on our test set. The results are presented in Table LABEL:Quantitative_result. We did not consider background noise when evaluating the experimental results. Due to the limited resolution of most previous work, all measurements are performed at a resolution of $512\times 512$. We report four metrics of different methods. FID gauges visual similarity, with lower values indicating better quality. PSNR measures distortion, and higher values imply less distortion. SSIM assesses luminance, contrast, and structure similarity, with values closer to 1 showing better structural preservation. LPIPS measures perceptual similarity and lower values mean closer resemblance to the reference. By comparing our model with these GAN-based and diffusion-based methods using the established metrics, our approach demonstrates a significant advantage over previous methods.
@@
Figure 8: Qualitative comparisons with existing methods. Given a line drawing and reference instances, our method demonstrates far more precise colorization and higher-quality results compared to other methods, effectively maintaining line-drawing structure and color consistency.
@@ 5.3 Ablation Study
Effectiveness of Color Matching. To assess our color-matching method, we remove the color matching component. This omission diminishes our model’s ability to retain color information and details from input references. For instance, in the second row of Figure 6, the colors of the bag and the character’s facial details are inconsistent.
@@
5.4 User Study
We developed a user interface to evaluate our proposed methods from user experiments. We invited 16 volunteers to experience our demonstration. Participants were required to perform reference-based colorization for all proposed models. The average testing time for each experimental set was more than 20 minutes. After the test was completed, we collected participants’ subjective evaluation scores in the following four dimensions: image quality, usability, reference similarity, and correlation between manipulated results. According to Figure 7, our method achieves the best performance in all four perspectives, especially in higher user satisfaction with image quality and control. Meanwhile, we can refine the proposed manipulation method further for better simplicity.
@@ 5.5 Limitations and Discussion
Our method has many advantages and potential applications but also faces several limitations. We summarize them as follows: (1) Flexible usage: When stylizing the same sketch with diverse reference images, our method retains the character’s identity. It adjusts details like lighting and background based on the reference styles. (2) Semantic awareness: For cartoon wallpapers or posters with multiple characters or objects, our approach uses multiple reference images to perform semantic-based colorization, ensuring each line art element gets colorized semantically. (3) Multi-object and occlusions: Single-subject line art can efficiently convert a sketch into a vivid, full-colored illustration, speeding up the production and enabling animators to test different color concepts quickly. However, in sketch images with many main objects or significant occlusions between them, the detailed inter-object colors may not be well maintained. (4) More explorations: The model can accurately color each element in multi-subject scenarios such as an anime battle scene or a forest scene while maintaining a harmonious overall color scheme.
@@ 6 Conclusion
This paper has presented MagicColor, a diffusion-based sketch colorization framework. Specifically, we introduce a multi-instance approach, a two-stage, self-play training strategy, and pixel-level color matching with edge loss. Our experiments demonstrated that MagicColor outperforms current methods in visual quality and style consistency, advancing the field of digital cartoon colorization. In future work, we plan to substantially augment the number of reference instances and enhance the model’s capacity to maintain semantic and color consistency. We also plan to release the source code for better collaboration between creative practitioners and AI researchers.

### v2

@@ 5.1 Implementation Detail
Dataset. The dataset employed in this study comprises two key components: the anime video dataset and the image dataset. The data were preprocessed from Sakuga [56], ATD-12K [66], and manually collected from the internet. A total of 1,000 image pairs were selected from animations to form the ablation test set, and the remaining data constituted the training set. We then extracted their instances as references.
@@
Hyper-parameters. During training, we resized the height and width of the input image to 512 and kept the original aspect ratio. Our model undergoes 100,000 optimization iterations with a batch size of 1. The learning rate is set to $1\times 10^{-5}$. The training phase takes around 7 days using 2 NVIDIA A800 80G GPUs. We perform a random horizon flip and random brightness adjustment for each reference image as data augmentation to simulate multi-view conditions.
@@ 5.2.1 Qualitative Results
To achieve the same settings, we use a complete reference image as the reference condition for the comparative baseline methods, and our method employs instances extracted from the same reference image. We compare our approach with previous sketch colorization methods. For GAN-based approaches, we consider RSIC [30] and SGA [36]. RSIC and SGA utilize GANs for colorization, each with a distinct architecture and training strategy. We observe that GAN-based methods tend to produce color incoordination. Moreover, when there are numerous color block regions in the image, color bleeding issues are likely to occur, where the color of one color block contaminates the adjacent color blocks. Regarding diffusion-based models, AnimeDiffusion [6], ColorizeDiffusion [89], and MangaNinja [44] (without point guided) perform better in terms of color control. However, as shown in Figure 5, they still struggle to effectively transfer the colors from the reference instances to the line drawing sketch, especially for the colors of smaller details in the images. The root cause is that the models fail to fully learn the color correspondences between the original image and the reference instances.
@@ 5.2.2 Quantitative Results
To comprehensively evaluate the colorization ability of our model, we quantitatively compared our method with the state-of-the-art colorization method on our test set. We constructed an Animation Dataset (1,570 pairs) from two anime films (Your Name and Spirited Away) and a Hand-Drawn Dataset (100 pairs) collected from real-user sketches as a test set. The results are presented in Table 3. We did not consider background noise when evaluating the experimental results. Due to the limited resolution of most previous work, all measurements are performed at a resolution of $512\times 512$. We report four metrics of different methods. FID gauges visual similarity, with lower values indicating better quality. PSNR measures distortion, and higher values imply less distortion. SSIM assesses luminance, contrast, and structure similarity, with values closer to 1 showing better structural preservation. LPIPS measures perceptual similarity, and lower values mean closer resemblance to the reference. By comparing our model with these GAN-based and diffusion-based methods using the established metrics, our approach demonstrates a significant advantage over previous methods.
@@ 5.3 Ablation Study
Effectiveness of Color Matching. Removing the color-matching component diminishes our model’s ability to retain color information and details from input references. For instance, in the second row of Figure 6, the colors of the bag and the character’s facial details are inconsistent.
@@ 5.4 Limitations and Discussion
Our method has many advantages and potential applications, but it also faces several limitations. We summarize them as follows: (1) Flexible usage: When stylizing the same sketch with diverse reference images, our method retains the character’s identity. It adjusts details like lighting and background based on the reference styles. (2) Semantic awareness: For cartoon wallpapers or posters with multiple characters or objects, our approach uses multiple reference images to perform semantic-based colorization, ensuring each line art element gets colorized semantically. (3) Multi-object and occlusions: Single-subject line art can efficiently convert a sketch into a vivid, full-colored illustration, speeding up the production and enabling animators to test different color concepts quickly. However, in sketch images with many main objects or significant occlusions between them, the detailed inter-object colors may not be well maintained. (4) More explorations: The model can accurately color each element in multi-subject scenarios such as an anime battle scene or a forest scene while maintaining a harmonious overall color scheme.
@@ 6 Conclusion
This paper has presented Follow-Your-Color, a diffusion-based sketch colorization framework. Specifically, we introduce a multi-instance approach, a two-stage, self-play training strategy, and pixel-level color matching with edge loss. Our experiments demonstrated that Follow-Your-Color outperforms current methods in visual quality and style consistency, advancing the field of digital cartoon colorization. In future work, we plan to substantially augment the number of reference instances and enhance the model’s capacity to maintain semantic and color consistency.
7 Acknowledagment
This work was partially supported by the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things #2023B1212010007. The Research Grant Council of the Hong Kong Special Administrative Region under grant number 16203122.